# Graph neural network outputs are almost surely asymptotically constant

### v1

@@ B.1 Combinatorial and growth lemmas
    $\left\lVert\frac{\sum_{v}\llbracket\rho(\bar{u},v)\rrbracket h(\llbracket\eta(\bar{u},v)\rrbracket)}{\sum_{v}h(\llbracket\eta(\bar{u},v)\rrbracket)}-\frac{f_{\pi}(\mathrm{H}(\bar{u}),t)}{g_{\pi}(\mathrm{H}(\bar{u}),t)}\right\rVert<\iota(\epsilon,\delta,\gamma)$
@@
    By Lemma B.6 it suffices to find $\zeta(\epsilon,\delta,\gamma),\xi(\epsilon,\delta,\gamma)>0$ which we can make arbitrarily small and constants $\nu,\Omega>0$ such that:
    $\left\lVert\frac{1}{n}\sum_{v}\llbracket\rho(\bar{u},v)\rrbracket h(\llbracket\eta(\bar{u},v)\rrbracket)-f_{\pi}(\mathrm{H}(\bar{u}),t)\right\rVert<\zeta(\epsilon,\delta,\gamma)$ (1)
    $\left\lVert\frac{1}{n}\sum_{v}h(\llbracket\eta(\bar{u},v)\rrbracket)-g_{\pi}(\mathrm{H}(\bar{u}),t)\right\rVert<\xi(\epsilon,\delta,\gamma)$ (2)
@@
    The proof of equation (1) and equation (2) is very similar. We prove equation (2) since it is slightly notationally lighter. Let:
    $g^{*}_{\pi}(\bar{u})\coloneqq\frac{1}{n}\sum_{v}h(e_{\eta}(\mathrm{H}(\bar{u},v),t))$
@@
    $\left\lVert g^{*}_{\pi}(\bar{u})-g_{\pi}(\mathrm{H}(\bar{u}),t)\right\rVert<\epsilon$
@@
    $\left(1-(\delta+\delta^{2})\right)\epsilon+(\delta+\delta^{2})2\kappa^{2}+\epsilon$
@@
    $\left\lVert e_{\rho}(\mathrm{H}(\bar{u},v),\mathrm{GrTp}(\bar{u},v))-\llbracket\rho(\bar{u},v)\rrbracket\right\rVert<\epsilon$ ($\dagger_{\rho}^{\textrm{l}oc}$)
@@
    $\left\lVert e_{\eta}(\mathrm{H}(\bar{u},v),\mathrm{GrTp}(\bar{u},v))-\llbracket\eta(\bar{u},v)\rrbracket\right\rVert<\epsilon$ ($\dagger_{\eta}^{\textrm{l}oc}$)
@@
    By Lemma B.3 there is $N_{\gamma}>N_{1}$ such that for all $n\geq N_{\gamma}$, with probability at least $1-\theta$, out of all the tuples $\bar{u}$ such that $\psi(\bar{u})\wedge\exists v\colon u_{i}Ev$, at least $1-\gamma$ satisfy that for at least $1-(\delta+\delta^{2})$ of the $v$ for which $\psi(\bar{u})\wedge u_{i}Ev$ holds also satisfy equation ($\dagger_{\rho}^{\textrm{l}oc}$) and equation ($\dagger_{\eta}^{\textrm{l}oc}$).
    Now, for any tuple $\bar{u}$ define the function:
    $f^{\circ}_{\pi}(v;\bar{u})\coloneqq e_{\rho}(\mathrm{H}(\bar{u},v),\mathrm{GrTp}(\bar{u},v))h(e_{\eta}(\mathrm{H}(\bar{u},v),\mathrm{GrTp}(\bar{u},v)))$
    Using a similar argument to the global weighted mean case, we find $N_{3}$ such that for all $n\geq N_{3}$ and all tuples $\bar{u}$ we have that:
    $\left\lVert f_{\pi}(\mathrm{H}(\bar{u}),\mathrm{GrTp}(\bar{u}))-\operatorname*{\mathbb{E}}_{v\in\operatorname{\mathcal{N}}(u_{i})}[f^{\circ}_{\pi}(v;\bar{u})]\right\rVert<\frac{1}{2}\epsilon$
    Notice that we now take the expectation with respect to the neighbourhood of $u_{i}$.
    Furthermore, taking $\Lambda$ as before, by Hoeffding’s Inequality (Theorem A.4) for any $\bar{u}$, the probability that:
    $\left\lVert\frac{1}{d(u_{i})}\sum_{v\in\operatorname{\mathcal{N}}(u_{i})}f^{\circ}_{\pi}(v;\bar{u})-\operatorname*{\mathbb{E}}_{v\in\operatorname{\mathcal{N}}(u_{i})}[f^{\circ}_{\pi}(v;\bar{u})]\right\rVert\geq\frac{1}{2}\epsilon$
@@
    $2d\exp\left(-\frac{\epsilon^{2}d(u_{i})}{2\Lambda^{2}}\right)$
    By Lemma B.2 we have that in all the non-sparse cases there is $R>0$ and $N_{4}$ such that for all $n\geq N_{4}$ with probability at least $1-\theta$, the proportion of all tuples $\bar{u}$ for which $\psi(\bar{u})$ holds such that:
    $d(u_{i})\geq R\log(n)$
    is at least $1-\delta$. For any such tuple $\bar{u}$ we have that:
    $2d\exp\left(-\frac{\epsilon^{2}d(u_{i})}{2\Lambda^{2}}\right)\leq 2dn^{-\frac{\epsilon^{2}R}{2\Lambda^{2}}}$
    There is $N_{5}\geq N_{4}$ such that for all $n\geq N_{5}$ we have:
    $2dn^{-\frac{\epsilon^{2}R}{2\Lambda^{2}}}<\theta\delta$
@@
    $\left\lVert\frac{1}{n}\sum_{v}f^{\circ}_{\pi}(v;\bar{u})-f_{\pi}(\mathrm{H}(\bar{u}),\mathrm{GrTp}(\bar{u}))\right\rVert<\epsilon$ ($\triangle_{f}^{\textrm{loc}}$)
@@
    $\left\lVert\frac{1}{n}\sum_{v}g^{\circ}_{\pi}(v;\bar{u})-g_{\pi}(\mathrm{H}(\bar{u}),\mathrm{GrTp}(\bar{u}))\right\rVert<\epsilon$ ($\triangle_{g}^{\textrm{loc}}$)
@@
To prove Theorem 5.1 for the non-sparse case, note that $e_{\tau}$ is only a function of the $\mathsf{GrTp}_{0}$ (since $\tau$ is a closed term), while $\mathsf{GrTp}_{0}$ consists of a single type, $\top$, which is satisfied by all graphs. So $e_{\tau}$ is a constant.
Appendix C Sparse Erdős-Rényi graphs
We now give the proof for the sparse case of Theorem 5.1. As discussed in the body, this requires us to analyze neighbourhood isomorphism types, which implicitly quantify over local neighbourhoods, rather than atomic types as in the non-sparse cases.
We need to define a notion of ‘particularisation’ for local isomorphism types, which is similar to the notion of type extension.
Definition C.1.
Let $\mathsf{IsoTp}_{k,\ell}$ be the set of all $(k,\ell)$ local isomorphism types.
Definition C.2.
Given a $(k,\ell)$ local isomorphism type $t=(T,\bar{u})$ and $u_{i}\in\bar{u}$, a particularisation of $t$ at $u_{i}$ is a $(k+1,\ell-1)$ local isomorphism type $t^{\prime}=(T^{\prime},\bar{u},v)$ where:
  1. (a)
    $v$ is a node in $T$ which is adjacent to $u_{i}$ and
  2. (b)
    $T^{\prime}$ is the subgraph of $T$ induced by $\operatorname{\mathcal{N}}_{\ell-1}^{T}(\bar{u},v)$.
Let $\operatorname{\mathrm{Par}}(t,u_{i})$ be the set of all particularisations of $t$ at $u_{i}$.
Definition C.3.
Given a $(k,\ell)$ local isomorphism type $t=(T,\bar{u})$ and $\ell^{\prime}\leq\ell$, the $\ell^{\prime}$-restriction of $t$, notation $t\restriction\ell^{\prime}$, is the $(k,\ell^{\prime})$ local isomorphism type $(\operatorname{\mathcal{N}}_{\ell^{\prime}}^{T}(\bar{u}),\bar{u})$.
For each $(k,\ell)$ local isomorphism type $t=(T,\bar{u})$, fix a canonical ordering of the nodes $T=\{s_{1},\ldots,s_{|t|}\}$ such that:
  • the first $|\bar{u}|$ nodes are $\bar{u}$, in order and
  • the canonical ordering of any $t\restriction\ell^{\prime}$ is an initial segment of the canonical ordering of $t$.
We will also assume that for every $(\operatorname{\mathcal{N}}_{\ell}(\bar{u}),\bar{u})$ we have a canonical ordering of $\operatorname{\mathcal{N}}_{\ell}(\bar{u})$ which agrees with the canonical ordering of the $(k,\ell)$ local isomorphism type of $\bar{u}$ (this is tantamount to choosing a canonical isomorphism from $(\operatorname{\mathcal{N}}_{\ell}(\bar{u}),\bar{u})$ onto the representative local isomorphism type with enumeration $\{s_{1},\ldots,s_{|t|}\}$).
Our first goal is to prove the Local Type Convergence Lemma, which we now recall:
See 5.7
For this we make use of the following bound on the neighbourhood sizes in the sparse case.
Lemma C.4.
Let $p$ satisfy the sparsity condition. Then for every $\ell\in\mathbb{N}$ and $\gamma>0$ there is $N>N_{2}$ such that for all $n\geq N$ with probability at least $1-\theta$ we have that for all nodes $u$:
$\frac{|\operatorname{\mathcal{N}}_{\ell}(u)|}{n}<\gamma$
@@ Appendix C Sparse Erdős-Rényi graphs
Let $p(n)=Kn^{-1}$. Take any node $u$. Then $d(u)$ is the sum of $n-1$ independent Bernoulli random variables with parameter $p(n)$. Hence by Chernoff’s Inequality (Theorem A.3):
@@
$\frac{|\operatorname{\mathcal{N}}_{\ell}(u)|}{n}<\frac{\log(n)^{\ell}}{n}$
@@
We can now prove the Local Type Convergence Lemma.
Proof of Lemma 5.7.
Let $p(n)=Kn^{-1}$. Take $k,\ell\in\mathbb{N}$ and $t\in\mathsf{IsoTp}_{k,\ell}$. First, by Lemma C.4, for every $\gamma>0$ there is $N_{\gamma}$ such that for all $n\geq N_{\gamma}$ with probability at least $1-\theta$ we have that for all nodes $u$:
$\frac{|\operatorname{\mathcal{N}}_{\ell}(u)|}{n}<\gamma$ ($\times$)
Let $q_{t}(n)$ be the probability that any given node $u$ satisfies $t$ (which is the same for every node). We can show inductively on $\ell$ that $q_{t}(n)$ is a polynomial in $n$ and $p(n)$. Note that every $n^{i}p(n)^{j}$ is monotone. Because $q_{t}(n)$ is also bounded, the two statements above imply that $q_{t}(n)$ converges to some limit $q_{t}$.
Now, let $P_{t}(n)$ be the proportion of nodes which satisfy $t$. We have that $\operatorname*{\mathbb{E}}[P_{t}(n)]=q_{t}(n)$. To complete the proof we need to show that $|P_{t}(n)-q_{t}(n)|$ can be made as small as we like. For this, we will bound the variance and use Chebyshev’s Inequality.
Given a node $v$, let $I_{t,v}$ be the indicator variable for the event that $v$ satisfies $t$. Then:
$P_{t}(n)=\frac{1}{n}\sum_{u\in V}I_{t,v}$
We have that:
$\mathrm{Var}(P_{t}(n))$ $=\frac{1}{n^{2}}\sum_{u,v\in V}\mathrm{Cov}(I_{t,v},I_{t,u})$
$=\frac{1}{n}\mathrm{Var}(I_{t,v})+\frac{n-1}{n}\mathrm{Cov}(I_{t,v},I_{t,u})$
where in the second line we fix some $u\neq v$, using the symmetry of the variances and covariances.
Since $I_{t,v}$ is a Bernoulli random variable, we have that:
$\mathrm{Var}(I_{t,v})=q_{t}(n)(1-q_{t}(n))<1$
Hence:
$\mathrm{Var}(P_{t}(n))$ $\leq\frac{1}{n}+\frac{n-1}{n}\mathrm{Cov}(I_{t,v},I_{t,u})$
$\leq\frac{1}{n}+\mathrm{Cov}(I_{t,v},I_{t,u})$ ($\ddagger$)
Turning to the covariance, we have that:
$\mathrm{Cov}(I_{t,v},I_{t,u})$ $=\operatorname*{\mathbb{E}}[I_{t,v}I_{t,u}]-\operatorname*{\mathbb{E}}[I_{t,v}]\operatorname*{\mathbb{E}}[I_{t,u}]$
$=\mathbb{P}(I_{t,v}=1\wedge I_{t,u}=1)-q_{t}(n)^{2}$
Using the Law of Total Probability, we have that:
$\mathbb{P}(I_{t,v}=1\wedge I_{t,u}=1)$
$\qquad=\left(\begin{array}[]{l}\mathbb{P}(I_{t,v}=1\wedge I_{t,u}=1\mid v\in\operatorname{\mathcal{N}}_{\ell}(u))\mathbb{P}(v\in\operatorname{\mathcal{N}}_{\ell}(u))\\ +\mathbb{P}(I_{t,v}=1\wedge I_{t,u}=1\mid v\notin\operatorname{\mathcal{N}}_{\ell}(u))\mathbb{P}(v\notin\operatorname{\mathcal{N}}_{\ell}(u))\end{array}\right)$
$\qquad\leq\mathbb{P}(v\in\operatorname{\mathcal{N}}_{\ell}(u))+\mathbb{P}(I_{t,v}=1\wedge I_{t,u}=1\mid v\notin\operatorname{\mathcal{N}}_{\ell}(u))$
Now, using the bound equation ($\times$×C) on $|\operatorname{\mathcal{N}}_{\ell}(u)|$ we have that:
$\mathbb{P}(v\in\operatorname{\mathcal{N}}_{\ell}(u))=\frac{|\operatorname{\mathcal{N}}_{\ell}(u)|-1}{n-1}<\frac{\gamma n-1}{n-1}\leq 2\gamma$
When $v\notin\operatorname{\mathcal{N}}_{\ell}(u)$, the distance between $u$ and $v$ is more than $\ell$. Since $t$ is local to the $\ell$-neighbourhood, in this case $I_{t,v}$ and $I_{t,v}$ are independent, so:
$\mathbb{P}(I_{t,v}=1\wedge I_{t,u}=1\mid v\notin\operatorname{\mathcal{N}}_{\ell}(u))=q_{t}(n)^{2}$
Hence:
$\mathrm{Cov}(I_{t,v},I_{t,u})\leq 2\gamma+q_{t}(n)^{2}-q_{t}(n)^{2}=2\gamma$
[…]
$\mathrm{Var}(P_{t}(n))\leq\frac{1}{n}+2\gamma$
By Chebyshev’s Inequality (Theorem A.2), for any $\epsilon>0$ we have:
$\mathbb{P}\left(|P_{t}(n)-q_{t}(n)|\geq\epsilon\right)\leq\frac{\mathrm{Var}(P_{t}(n))}{\epsilon^{2}}\leq\frac{\frac{1}{n}+2\gamma}{\epsilon^{2}}$
Since we have control over $\gamma$, by making $N_{\gamma}$ large enough we can make this probability as small as we like for large enough $n$. Thus, the proportion of nodes which satisfy $t$ converges to $q_{t}$. ∎
In the proof of Theorem 5.1 we will also make use of the following corollary to the Local Type Convergence Lemma, which concerns the convergence of types for single-node extensions of tuples.
Corollary C.5.
Let $p$ satisfy sparsity. For every $k,\ell\in\mathbb{N}$ and $t\in\mathsf{IsoTp}_{k+1,\ell}$ there is $q_{t}\in[0,1]$ such that for all $\epsilon,\theta>0$ there is $N\in\mathbb{N}$ such that for all $n\geq N$ with probability at least $1-\theta$ we have that for all $k$-tuples of nodes $\bar{u}$:
||$\left|\frac{|\{v\mid\mathrm{IsoTp}_{\ell}(\bar{u},v)=t\}|}{n}-q_{t}\right|<\epsilon$
Proof.
Take $k,\ell\in\mathbb{N}$ and $t\in\mathsf{IsoTp}_{k+1,\ell}$. There are two cases.
@@
There is a path from $v$ to some node $u_{i}$ of $\bar{u}$ in $t$.
@@
$\{v\mid\mathrm{IsoTp}_{\ell}(\bar{u},v)=t\}\subseteq\operatorname{\mathcal{N}}_{\ell}(u_{i})$
By Lemma C.4, for every $\epsilon>0$ there is $N$ such that for all $n\geq N$ with probability at least $1-\theta$ we have that for all nodes $v$:
$\frac{|\operatorname{\mathcal{N}}_{\ell}(v)|}{n}<\epsilon$
@@
||$\left|\frac{|\{v\mid\mathrm{IsoTp}_{\ell}(\bar{u},v)=t\}|}{n}\right|<\left|\frac{|\operatorname{\mathcal{N}}_{\ell}(u_{i})|}{n}\right|<\epsilon$
@@
There is no path from $v$ to any node of $\bar{u}$ in $t$.
Let $t^{*}$ be the $(1,\ell)$ local isomorphism type of $v$ in $t$. Then for all tuples $\bar{u}$:
$\{v\mid\mathrm{IsoTp}_{\ell}(\bar{u},v)=t\}=\{v\mid\mathrm{IsoTp}_{\ell}(v)=t^{*}\}\setminus\operatorname{\mathcal{N}}_{\ell}(\bar{u})$
@@
|$\frac{|\{v\mid\mathrm{IsoTp}_{\ell}(v)=t^{*}\}|}{n}-\frac{|\operatorname{\mathcal{N}}_{\ell}(\bar{u})|}{n}\leq\frac{|\{v\mid\mathrm{IsoTp}_{\ell}(\bar{u},v)=t\}|}{n}$
@@
|$\frac{|\{v\mid\mathrm{IsoTp}_{\ell}(\bar{u},v)=t\}|}{n}\leq\frac{|\{v\mid\mathrm{IsoTp}_{\ell}(v)=t^{*}\}|}{n}$

### v2

@@ C.1 Combinatorial and growth lemmas
    $\left\lVert\frac{\sum_{v}\llbracket\rho(\bar{u},v)\rrbracket h(\llbracket\eta(\bar{u},v)\rrbracket)}{\sum_{v}h(\llbracket\eta(\bar{u},v)\rrbracket)}-\frac{f_{\pi}(H_{G}(\bar{u}),t)}{g_{\pi}(H_{G}(\bar{u}),t)}\right\rVert<\iota(\epsilon,\delta,\gamma)$
@@
    By Lemma C.7 it suffices to find $\zeta(\epsilon,\delta,\gamma),\xi(\epsilon,\delta,\gamma)>0$ which we can make arbitrarily small and constants $\nu,\Omega>0$ such that:
    $\left\lVert\frac{1}{n}\sum_{v}\llbracket\rho(\bar{u},v)\rrbracket h(\llbracket\eta(\bar{u},v)\rrbracket)-f_{\pi}(H_{G}(\bar{u}),t)\right\rVert<\zeta(\epsilon,\delta,\gamma)$ (1)
    $\left\lVert\frac{1}{n}\sum_{v}h(\llbracket\eta(\bar{u},v)\rrbracket)-g_{\pi}(H_{G}(\bar{u}),t)\right\rVert<\xi(\epsilon,\delta,\gamma)$ (2)
@@
    The proofs of equation (1) and equation (2) are very similar. We prove equation (2) since it is slightly notationally lighter. Let:
    $g^{*}_{\pi}(\bar{u})\coloneqq\frac{1}{n}\sum_{v}h(e_{\eta}(H_{G}(\bar{u},v),t))$
@@
    $\left\lVert g^{*}_{\pi}(\bar{u})-g_{\pi}(H_{G}(\bar{u}),t)\right\rVert<\epsilon$
@@
    $\left\lVert g_{\pi}(H_{G}(\bar{u}),t)-\frac{1}{n}\sum_{v}h(\llbracket\eta(\bar{u},v)\rrbracket)\right\rVert<\left(1-(\delta+\delta^{2})\right)\epsilon+(\delta+\delta^{2})2\kappa^{2}+\epsilon$
@@
    $\left\lVert e_{\rho}(H_{G}(\bar{u},v),\mathrm{GrTp}(\bar{u},v))-\llbracket\rho(\bar{u},v)\rrbracket\right\rVert<\epsilon$ ($\dagger_{\rho}^{\textrm{l}oc}$)
@@
    $\left\lVert e_{\eta}(H_{G}(\bar{u},v),\mathrm{GrTp}(\bar{u},v))-\llbracket\eta(\bar{u},v)\rrbracket\right\rVert<\epsilon$ ($\dagger_{\eta}^{\textrm{l}oc}$)
@@
    By Lemma C.4 there is $N_{\gamma}>N_{1}$ such that for all $n\geq N_{\gamma}$, with probability at least $1-\theta$, out of all the tuples $\bar{u}$ such that $\psi(\bar{u})\wedge\exists v\colon u_{i}Ev$, at least $1-\gamma$ satisfy that for at least $1-(\delta+\delta^{2})$ of the $v$ for which $\psi(\bar{u})\wedge u_{i}Ev$ holds also satisfy equation ($\dagger_{\rho}^{\textrm{l}oc}$) and equation ($\dagger_{\eta}^{\textrm{l}oc}$).
    Now consider $t(\bar{u})\in\mathsf{GrTp}_{k}$ and take any $t^{\prime}(\bar{u},v)\in\mathsf{Ext}_{u_{i}}(t)$. For any tuple $\bar{u}$ of nodes such that $\mathrm{GrTp}(\bar{u})=t$, define:
    $\llbracket t^{\prime}(\bar{u})\rrbracket\coloneqq\{v\in\operatorname{\mathcal{N}}(u_{i})\mid t^{\prime}(\bar{u},v)\}$
    By Lemma C.3 we have that in all the non-sparse cases there is $R>0$ and $N_{3}$ such that for all $n\geq N_{3}$ with probability at least $1-\theta$, the proportion of all tuples $\bar{u}$ for which $\psi(\bar{u})$ holds such that:
    $d(u_{i})\geq R\log(n)$
    is at least $1-\delta$. Then, as before, there is $N_{4}\geq N_{3}$ such that for all $n\geq N_{4}$, probability at least $1-\theta$, for all pairs $(t,t^{\prime})$ the proportion of all tuples $\bar{u}$ for which $\psi(\bar{u})$ and $t(\bar{u})$ hold such that:
    $\left|\frac{|\llbracket t^{\prime}(\bar{u})\rrbracket|}{d(u_{i})}-\alpha(t,t^{\prime})\right|<\frac{1}{2^{k+1}}\epsilon$
    For any tuple $\bar{u}$ define the function:
    $f^{\circ}_{\pi}(\bar{u},v)\coloneqq e_{\rho}(H_{G}(\bar{u},v),\mathrm{GrTp}(\bar{u},v))h(e_{\eta}(H_{G}(\bar{u},v),\mathrm{GrTp}(\bar{u},v)))$
    Taking $\Lambda$ as before, by Hoeffding’s Inequality for any $\bar{u}$ the probability that:
    $\left\lVert\frac{1}{d(u_{i})}\sum_{v\in\operatorname{\mathcal{N}}(u_{i})\mid t^{\prime}(\bar{u},v)}f^{\circ}_{\pi}(\bar{u},v)-\frac{|\llbracket t^{\prime}(\bar{u})\rrbracket|}{d(u_{i})}\operatorname*{\mathbb{E}}_{{\bm{b}}\sim\mathbb{F}}\left[e_{\rho}(H_{G}(\bar{u},{\bm{b}}),t^{\prime})h(e_{\eta}(H_{G}(\bar{u},{\bm{b}}),t^{\prime}))\right]\right\rVert$
    $\geq\frac{1}{2^{k+1}}\epsilon$
@@
    $2d\exp\left(-\frac{\epsilon^{2}|\llbracket t^{\prime}(\bar{u})\rrbracket|}{2\Lambda^{2}}\right)$
    Using a similar argument to the global weighted mean case, there is $N_{5}$ such that for all $n\geq N_{5}$, with probability at least $1-\theta$ for all $t^{\prime}$ such that $\alpha(t,t^{\prime})>0$ we have that for every at least $1-\delta$ of the tuples $\bar{u}$ such that $\psi(\bar{u})$ and $t(\bar{u})$ hold:
    $2d\exp\left(-\frac{\epsilon^{2}|\llbracket t^{\prime}(\bar{u})\rrbracket|}{2\Lambda^{2}}\right)<\theta\delta$
@@
    $\left\lVert\frac{1}{n}\sum_{v}f^{\circ}_{\pi}(\bar{u},v)-f_{\pi}(H_{G}(\bar{u}),\mathrm{GrTp}(\bar{u}))\right\rVert<\epsilon$ ($\triangle_{f}^{\textrm{loc}}$)
@@
    $\left\lVert\frac{1}{n}\sum_{v}g^{\circ}_{\pi}(\bar{u},v)-g_{\pi}(H_{G}(\bar{u}),\mathrm{GrTp}(\bar{u}))\right\rVert<\epsilon$ ($\triangle_{g}^{\textrm{loc}}$)
@@
To prove Theorem C.1, note that $e_{\tau}$ is only a function of the $\mathsf{GrTp}_{0}$ (since $\tau$ is a closed term), while $\mathsf{GrTp}_{0}$ consists of a single type, $\top$, which is satisfied by all graphs. So $e_{\tau}$ is a constant.
Appendix D Sparse Erdős-Rényi and Barabási-Albert
[…]
Theorem D.1.
Consider $(\mu_{n})_{n\in\mathbb{N}}$ sampling a graph $G$ from either of the following models and node features independently from i.i.d. bounded distributions on $d$ features.
  1. 1.
    The Erdős-Rényi distribution $\mathrm{ER}(n,p(n))$ where $p$ satisfies Sparsity: for some $K>0$ we have: $p(n)=Kn^{-1}$.
  2. 2.
    The Barabási-Albert model $\mathrm{BA}(n,m)$ for any $m\geq 1$.
[…]
As discussed in the body, this requires us to analyze neighbourhood isomorphism types, which implicitly quantify over local neighbourhoods, rather than atomic types as in the non-sparse cases.
Remark D.2.
Formally, a $k$-rooted graph is a tuple $(T,\bar{u})$. At various points below to lighten the notation we drop the ‘$\bar{u}$’ and refer simply to a $k$-rooted graph $T$.
D.1 Combinatorial tools for the sparse case
As in the dense case, before we turn to analysis of our term language, we prove some results about the underlying random graph model that we can use as tools.
The key combinatorial tool we will use is the following, which states that for any tuple of nodes $\bar{u}$, the proportion of nodes $v$ such that $(\bar{u},v)$ has any particular neighbourhood type converges.
Theorem D.3.
Consider sampling a graph $G$ from either the sparse Erdős-Rényi or Barabási-Albert distributions. Let $(T,(\bar{w},y))$ be a $(k+1)$-rooted graph and $\ell\in\mathbb{N}$. There is $q_{T}\in[0,1]$ such that for all $\epsilon,\theta>0$ there is $N\in\mathbb{N}$ such that for all $n\geq N$ with probability at least $1-\theta$ we have that for all $k$-tuples of nodes $\bar{u}$:
||$\left|\frac{|\{v\mid\operatorname{\mathcal{N}}^{G}_{\ell}(\bar{u},v)\cong(T,(\bar{w},y))\}|}{n}-q_{T}\right|<\epsilon$
This follows from two important properties shared by these distributions. The first is what’s known as “weak local convergence”, which is essentially a non-parametrised version of Theorem D.3.
Definition D.4.
A sequence $(\mu_{n})_{n\in\mathbb{N}}$ of graph distributions has weak local convergence if for every $k$-rooted graph $(T,\bar{w})$ and $\ell\in\mathbb{N}$ there is $q_{T}\in[0,1]$ such that for all $\epsilon,\theta>0$ there is $N\in\mathbb{N}$ such that for all $n\geq N$ with probability at least $1-\theta$ we have that:
||$\left|\frac{|\{\bar{u}\mid\operatorname{\mathcal{N}}^{G}_{\ell}(\bar{u})\cong(T,\bar{w})\}|}{n}-q_{T}\right|<\epsilon$
Theorem D.5.
The sparse Erdős-Rényi and Barabási-Albert distributions have weak local convergence.
@@ D.1 Combinatorial tools for the sparse case
See [51, Theorem 2.18] for the sparse Erdős-Rényi distribution and [17, Theorem 4.2.1] for the Barabási-Albert distribution. ∎
The second property is that the maximum neighbourhood sizes of nodes grows slower than $n$.
Lemma D.6.
Consider sampling a graph $G$ from either the sparse Erdős-Rényi or Barabási-Albert distributions. Then for every $\ell\in\mathbb{N}$ and $\gamma>0$ there is $N>N_{2}$ such that for all $n\geq N$ with probability at least $1-\theta$ we have that for all nodes $u$:
$\frac{|\operatorname{\mathcal{N}}^{G}_{\ell}(u)|}{n}<\gamma$
Proof.
For the Barabási-Albert distribution we appeal to a well-known result that if $M_{n}$ is the maximum degree of the $n$-vertex Barabási-Albert graph then for every $\epsilon>0$ we have that:
$\mathbb{P}(M_{n}<n^{\frac{1}{2}+\epsilon})\to 1\text{ as }n\to\infty$
See [41, 38, 39].
Now consider the sparse Erdős-Rényi case. Let $p(n)=Kn^{-1}$. Take any node $u$. Then $d(u)$ is the sum of $n-1$ independent Bernoulli random variables with parameter $p(n)$. Hence by Chernoff’s Inequality (Theorem B.3):
@@
$\frac{|\operatorname{\mathcal{N}}^{G}_{\ell}(u)|}{n}<\frac{\log(n)^{\ell}}{n}$
@@
With these two properties we can prove our combinatorial result.
Proof of Theorem D.3.
There are two cases.
@@
There is a path from $y$ to some node $w_{i}$ of $\bar{w}$ in $T$.
@@
$\{v\mid\operatorname{\mathcal{N}}^{G}_{\ell}(\bar{u},v)\cong(T,\bar{w},y)\}\subseteq\operatorname{\mathcal{N}}^{G}_{\ell}(u_{i})$
By Lemma D.6, for every $\epsilon>0$ there is $N$ such that for all $n\geq N$ with probability at least $1-\theta$ we have that for all nodes $v$:
$\frac{|\operatorname{\mathcal{N}}^{G}_{\ell}(v)|}{n}<\epsilon$
@@
||$\left|\frac{|\{v\mid\operatorname{\mathcal{N}}^{G}_{\ell}(\bar{u},v)\cong(T,\bar{w},y)\}|}{n}\right|<\left|\frac{|\operatorname{\mathcal{N}}^{G}_{\ell}(u_{i})|}{n}\right|<\epsilon$
@@
There is no path from $y$ to any node of $\bar{w}$ in $T$.
Then for all tuples $\bar{u}$:
$\{v\mid\operatorname{\mathcal{N}}^{G}_{\ell}(\bar{u},v)\cong(T,\bar{w},y)\}=\{v\mid\operatorname{\mathcal{N}}^{G}_{\ell}(v)\cong\operatorname{\mathcal{N}}^{T}_{\ell}(y)\}\setminus\operatorname{\mathcal{N}}^{G}_{\ell}(\bar{u})$
@@
|$\frac{|\{v\mid\operatorname{\mathcal{N}}^{G}_{\ell}(v)\cong\operatorname{\mathcal{N}}^{T}_{\ell}(y)\}|}{n}-\frac{|\operatorname{\mathcal{N}}^{G}_{\ell}(\bar{u})|}{n}\leq\frac{|\{v\mid\operatorname{\mathcal{N}}^{G}_{\ell}(\bar{u},v)\cong(T,\bar{w},y)\}|}{n}$
@@
|$\frac{|\{v\mid\operatorname{\mathcal{N}}^{G}_{\ell}(\bar{u},v)\cong(T,\bar{w},y)\}|}{n}\leq\frac{|\{v\mid\operatorname{\mathcal{N}}^{G}_{\ell}(v)\cong\operatorname{\mathcal{N}}^{T}_{\ell}(y)\}|}{n}$